OpenStudy (precal):

A man 6 feet talk walks at a rate of 5 per second toward a street light that is 16 feet tall. At what rate is the length of his shadow changing when he is 10 feet from the base of the light?

OpenStudy (anonymous):

 \(x\) is the length of the man's shadow, \(h\) is the distance between the man and the light. You're given that \(\dfrac{dh}{dt}=-5\) (-5 because \(h\) is decreasing), and you have to find \(\dfrac{dx}{dt}\). Similar triangles give the ratio, \[\frac{6}{x}=\frac{16}{x+h}\]